Bill and Will are brothers who are both taking graduate level art classes. To help pay off their student loans, they decide to sell watercolor replicas of Andy Warhol’s art. To produce one piece of work, Bill takes 0.25 hours less than Will. Working together, it takes them 3 hours. Find the time it would take each of them if they worked individually. Round your answers to 2 decimal places.
Answer by stanbon(75887)

To produce one piece of work, Bill takes 0.25 hours less than Will. Working together, it takes them 3 hours. Find the time it would take each of them if they worked individually. Round your answers to 2 decimal places.
--------
Will's rate = 1/x job/hr
Bill's rate = 1/(x-0.25) job/hr
Together rate = 1/3 job/hr
-------
Equation
rate + rate = together rate
-------
1/x + 1/(x-0.25) = 1/3
--------
3(x-0.25) + 3x = x(x-0.25)
-------
3x - 0.75 + 3x = x^2 - 0.25x
--------
x^2 - 6.25x + 0.75 = 0
--------
x = 6.13 hrs (Will's time to do the job alone)
x-0.25 = 5.88 hrs (Bill's time to do the job alone)
============
Cheers,
Stan H.